n a certain community, levels of air pollution may exceed federal standards for ozone or for particulate matter on some days. For a particular summer season, let X be the number of days on which the ozone standard is exceeded and let Y be the number of days on which the particulate matter standard is exceeded. Assume that the joint probability mass function of X and Y is given in the following table:

y

x

0

1

2

0

0.10

0.11

0.05

1

0.17

0.23

0.08

2

0.06

0.14

0.06

Find the conditional probability mass function of pY|X(y|0). Round the answers to four decimal places.

The value of pY|X(0|0) is .

The value of pY|X(1|0) is .

The value of pY|X(2|0) is .
Answer #1

P(X = 0) = 0.1 + 0.11 + 0.05 = 0.26

P(Y = 0 = | X = 0) = P(Y = 0 and X = 0) / P(X = 0) = 0.1 / 0.26 = 0.3846

P(Y = 1 = | X = 0) = P(Y = 1 and X = 0) / P(X = 0) = 0.11 / 0.26 = 0.4231

P(Y = 2 = | X = 0) = P(Y = 2 and X = 0) / P(X = 0) = 0.05 / 0.26 = 0.1923